A car travelling at a speed of 50 mph. The driver saw a roadblock 80m ahead and stepped on the brake causing the car to decelerate uniformly at 10 m/s². Find the distance from the roadblock to the point where the car stopped. Assume perception reaction time is 2 seconds.


1
Expert's answer
2021-10-18T11:02:03-0400

"v=50~\frac{mi}h=22.35~\frac ms,"

"l'=l-vt=35.3~m,"

"s=l'-\frac{v^2}{2a}=10.3~m."